OpenStudy (anonymous):

Rachel drove 845 miles in 13 hours. At the same rate, how long would it take her to drive 585 miles?

OpenStudy (anonymous):

Use your ratios recall cross multiplication just have the same units on top!\[\frac{ 13 hr }{ 845 } = \frac{ x }{ 585 }\] can you solve for x?

OpenStudy (anonymous):

Yes thank you!

OpenStudy (anonymous):

Wait no I can't.

OpenStudy (anonymous):

lol ok, alright here try this \[x = \frac{ 13 \times 585 }{ 845 }\]

OpenStudy (anonymous):

Ok, I think I get it now, Thank you!

OpenStudy (anonymous):

yw =))